OpenStudy (anonymous):

Which of the following is the slope-intercept form of the equation -6x+y=9?

OpenStudy (anonymous):

A.y=-6x-9 B.y=-6x+9 C.y=6x+9 D.y=6x-9

OpenStudy (anonymous):

y = 6x + 9

OpenStudy (anonymous):

thank ya :)